PLEASE HELP ME IF ANY ONE CAN..........
IF X^2/3=X,THEN X=
(A)0 OR -3
(B)3 OR -3
(C) 3 ONLY
(D)0 OR 3
PLEASE ANSWER ACCORDING TO THE CAPS

Answer by ewatrrr(24785)

x^2 = 3x
x^2 - 3x = 0
x(x-3) = 0
......
x = 0
0r
(x-3) = 0, x = 3
(D)x = 0 or 3